# O-RAN xApps Conflict Management using Graph Convolutional Networks

### v1

@@ Abstract
Open Radio Access Network (O-RAN) adopts a flexible, open, and virtualized structure with standardized interfaces, reducing dependency on a single supplier. Conflict management in O-RAN refers to the process of identifying and resolving conflicts between network applications. xApps are applications deployed at the RAN Intelligent Controller (RIC) that leverage advanced AI/ML algorithms to make dynamic decisions for network optimization. The lack of a unified mechanism to coordinate and prioritize the actions of different applications can create three types of conflicts (direct, indirect, and implicit). In our paper, we introduce a novel data-driven GCN-based method called Graph-based xApps Conflict and Root Cause Analysis Engine (GRACE) based on Graph Convolutional Network (GCN). It detects three types of conflicts (direct, indirect, and implicit) and pinpoints the root causes (xApps). GRACE captures the complex and hidden dependencies among the xApps, the controlled parameters, and the KPIs in O-RAN to detect possible conflicts. Then, it identifies the root causes (xApps) contributing to the detected conflicts. The proposed method was tested on highly imbalanced datasets where the number of conflict instances ranges from 40% to 10%. The model is tested in a setting that simulates real-world scenarios where conflicts are rare to assess its performance and generalizability. Experimental results demonstrate an exceptional performance, achieving a high F1-score greater than 98% for all the case studies.
@@
5G Networks and Beyond, O-RAN, Quality of Service, Conflict Management, xApps, Graph Neural Network, Root Cause Analysis, Key Performance Indicators.
@@ I Introduction
Other studies seek to detect or predict the occurrences of conflicts in O-RAN. In [16], a team deep Q-learning algorithm for resource allocation is disclosed to mitigate xApps conflicts in O-RAN. In [17], a bargaining game-theoretic approach is introduced along with a standardized Conflict Mitigation System (CMS) and independent controllers. This work is further improved in [18]. They use a cooperative game theory to optimize the conflicting parameters while maximizing the number of xApps that meet QoS requirements. In [19], a Deep Q-Network and a multi-headed multilayer perceptron network model are presented. The method distills knowledge from multiple xApps to train a single model retaining the capabilities of previous xApps.
Other works address the xApps conflict management using graph-based approaches. The authors in [20] aims to detect, characterize, and mitigate conflicts among applications in O-RAN. Accordingly, they combine hierarchical graphs with statistical knowledge of applications to find dependencies between control parameters and Key Performance Measurements. In [21], the paper suggests using GraphSAGE for reconstructing and labeling conflict graphs in O-RAN. It captures the hidden dependencies between xApps, parameters, and KPIs. Then, graph labeling is performed to identify the different types of conflicts based on well-defined, graph-based definitions.
Understanding the complex and intricate dependencies between the xApps, the controlled parameters, and the KPIs is imperative to address the xApp conflict RCA. It allows us to detect the different types of conflicts in real time and derive the contributing root causes (xApps) for proactive maintenance. In the previously mentioned non-graph-based approaches, there is still a substantial limitation in their capacity to effectively capture and model dependencies within complex data structures. The proposed method in [21] leverages GraphSAGE to effectively capture the dependencies within the data. However, it still falls short of fully accounting for the imbalanced nature of the datasets in practical life scenarios, as these conflicts are rare to happen. Addressing this limitation is crucial for enhancing the robustness and generalizability of the proposed approach in real-world applications. Accordingly, we propose GRACE, a GCN-based method that is capable of handling highly imbalanced datasets with up to 10% conflicts. It detects the xApps conflicts (direct, indirect, and implicit) and identifies their root causes (xApps) for future mitigation strategies.
@@
  • We propose a 5G O-RAN xApps conflict detection with a root cause analysis method using Graph Convolutional Networks (GCNs). This allows us to capture the complex and hidden dependencies between the xApps, parameters, and KPIs in the network to detect xApps conflicts (direct, indirect, and implicit).
  • We leverage a binary state-based GCN structure to map and explain the complex relationships. This ensures the generalizability of the proposed solution, which can be extended to other networks. This structure along with the detections are the key inputs for the root cause analysis.
  • We use the focal loss function in the model training. This allows the model to handle highly imbalanced datasets. The model generalizes effectively to unseen data while maintaining an F1 score greater than 98%.
The remainder of this paper is structured as follows. Section II presents the problem formulation, clearly defining the research question and outlining the scope and objectives of the study. Section III describes the proposed system model. Section IV details the evaluation metrics, the experimental setup, and the interpretation of the results. Section V concludes this paper.
@@ II Preliminaries
Figure 1: The xApps conflicts examples.
@@ III Proposed Method
A database is required to store the collected data from the SM in the Near-RT RIC using the messaging infrastructure. Then, the proposed system retrieves the collected data from the database for further processing using a shared data layer. Our proposed solution requires the implementation of a component that converts the collected data into a binary-state dataset as depicted in Fig. 4. In particular, the state of an activated xApp is presented by ”1”, and a deactivated xApp is represented by ”0”. Similarly, if the value of a parameter or a KPI is changed from its previous timestamp value, the state is set to ”1”, otherwise, the state is set to ”0”.
The proposed solution executes three major tasks, namely Graph Structure Creator (GSC), Graph Anomaly Detector (GAD), and Root Cause Analyst (RCA). The high-level diagram of the solution is illustrated in Fig. 3. In the remainder of this section, we will provide a detailed description of each module in the proposed method, outlining the corresponding steps and functionality.
@@
Figure 3: The high-level diagram of GRACE.
III-A Dataset Generation and Conflicts Modeling
@@ III-B Graph Structure Creator (GSC)
GSC module is responsible for creating graph-structured data from the binary-state dataset comprising the collected data from the Near-RT RIC (see Fig. 5). A graph-structured data is needed as the GCN-based model operates on graph data. It ingests the binary-state dataset and the number of xApps, parameters, and KPIs as input. The resultant graph-structured data will be the combination of three subgraphs ($G^{\mathrm{PA}}$, $G^{\mathrm{KP}}$, and $G^{\mathrm{P^{\prime}P}}$).
Figure 5: The graph structure creator module steps.
@@
$G^{\mathrm{PA}}$ subgraph: This subgraph $G^{\mathrm{PA}}=\left(V^{\mathrm{PA}},E^{\mathrm{PA}}\right)$ represents the relationships between the applications and the control parameters. Nodes of $G^{\mathrm{PA}}$ are both the applications and the control parameters, i.e., $V^{\mathrm{PA}}=\mathcal{P}\cup\mathcal{A}$, and edges $E^{\mathrm{PA}}$ represent whether or not an application is controlling a parameter $p\in\mathcal{P}$. Any 2-tuple $(p,a)\in\mathcal{P}\times\mathcal{A}$ is an edge of $G^{\mathrm{PA}}$, i.e., $(p,a)\in E^{\mathrm{PA}}$ if and only if parameter $p$ is directly controlled by an application $a$. This will cause a direct conflict.
@@
$G^{\mathrm{KP}}$ subgraph: This subgraph $G^{\mathrm{KP}}=\left(V^{\mathrm{KP}},E^{\mathrm{KP}}\right)$ represents the relationships between the control parameters and the KPIs. Nodes of $G^{\mathrm{KP}}$ are both the parameters and the KPIs, i.e., $V^{\mathrm{KP}}=\mathcal{K}\cup\mathcal{P}$, and edges $E^{\mathrm{KP}}$ represent whether or not a parameter is modifying the value of a KPI $p\in\mathcal{P}$. Any 2-tuple $(k,p)\in\mathcal{K}\times\mathcal{P}$ is an edge of $G^{\mathrm{KP}}$, i.e., $(k,p)\in E^{\mathrm{KP}}$ if and only if KPI $k$ is directly affected by a parameter $p$. This will cause an indirect conflict.
@@
$G^{\mathrm{P^{\prime}P}}$ subgraph: This subgraph $G^{\mathrm{P^{\prime}P}}=\left(V^{\mathrm{P^{\prime}P}},E^{\mathrm{P^{\prime}P}}\right)$ represents the relationships between the control parameters themselves. Nodes of $G^{\mathrm{P^{\prime}P}}$ are both the parameters in the sets $\mathcal{P^{\prime}}$ and $\mathcal{P}$, i.e., $V^{\mathrm{P^{\prime}P}}=\mathcal{P^{\prime}}\cup\mathcal{P}$, and edges $E^{\mathrm{P^{\prime}P}}$ represent whether or not a parameter $p\in\mathcal{P}$ is modifying the value of another parameter $p^{\prime}\in\mathcal{P^{\prime}}$. Any 2-tuple $(p^{\prime},p)\in\mathcal{P^{\prime}}\times\mathcal{P}$ is an edge of $G^{\mathrm{P^{\prime}P}}$, i.e., $(p^{\prime},p)\in E^{\mathrm{P^{\prime}P}}$ if and only if a parameter $p^{\prime}$ is directly affected by a parameter $p$. This will cause an implicit conflict.
@@
To construct the graph $G^{\mathrm{PA}}$, the GSL iterates over the applications and the parameters nodes and stores their corresponding state values. Subsequently, the resultant graph nodes are xApps and the parameters. The edges represent the nodes having a state ”1” concurrently. The edges reflect the nodes that are affecting each other at the same time. This leads to direct conflicts.
@@
Similarly, to build the graph $G^{\mathrm{KP}}$, the GSL iterates over the parameters and the KPIs nodes and stores their corresponding state values. The output graph nodes are the parameters and the KPIs. The edges indicate the parameters and KPIs whose values change simultaneously by having a state ”1, resulting in implicit conflicts.
@@
Lastly, to create the graph $G^{\mathrm{P^{\prime}P}}$, the GSL iterates over the parameters and checks their state values. The nodes of the graphs are the parameters, and the edges refer to the links among the parameters that have a state ”1”. This means the parameters are impacting each other behavior. This produces indirect conflicts.
@@
During the construction of the three subgraphs, unique node features and edge attributes are assigned to the nodes and their edges that are contributing to the same type of conflict. GCN uses node features as the starting point to update the node embeddings (representations) through multiple layers of convolutions. These embeddings are refined iteratively as the model propagates information across neighboring nodes. Including edge attributes in the graph convolution process allows the model to learn a more nuanced understanding of node interactions. When performing convolution, edge attributes help the GCN understand the nature of the relationship between nodes.
III-C Graph Anomaly Detector (GAD)
@@
The GAD module is responsible for detecting whether the instance is normal or whether we have a conflict (see Fig. 6). The output of GAD will be one of four labels (0: normal, 1: direct conflict, 2: implicit conflict, and 3: indirect conflict). The fully constructed graph structure from the GSL module is fed to the GAD module. It consists of a lightweight Graph Convolutional Network (GCN). GCN learns node-level representations from the graph-structured data, which are then aggregated via a global mean pooling into a graph-level representation for final classification. Then, a fully connected neural network (NN) layer is used to learn the final transformation from the graph’s representation to the output classes.
Figure 6: The graph anomaly detector module steps.
@@
To address the imbalances in the datasets, the classifier is enhanced by a focal loss function [24]. It is an enhancement to the standard cross-entropy loss by adding two key parameters, alpha and gamma. This function adjusts the loss contribution of easy and hard examples. It focuses more on hard-to-classify examples that are misclassified by the model. In parallel, it down-weights the loss associated with easy-to-classify examples. Alpha adjusts for class imbalance by balancing the weight between different classes. Gamma focuses the model on harder examples by reducing the loss for easy, well-classified examples. This helps to prevent the model from being overwhelmed by the majority class in imbalanced datasets and gives more attention to challenging instances. We compute alpha using inverse class weights by calculating the inverse frequency of each class in the dataset. On the other hand, gamma is manually tuned for model optimization. The focal loss function for multi-class classification is given by:
@@
Instead of processing the entire graph at once, mini-batch processing is executed. This involves processing smaller subgraphs simultaneously, which helps reduce memory usage and computational time. Furthermore, a stratified K-Fold is implemented to ensure that each fold in the cross-validation process has a similar distribution of classes as the entire dataset during training and validation. This is particularly useful since our datasets are highly imbalanced. L2 regularization (also known as Ridge regularization) [25] is implemented by tuning the weight decay parameter. This helps to mitigate overfitting by adding a penalty to the loss function that prevents the model from giving too much weight to any single feature. It achieves this by squaring the model’s weights and adding the result to the loss, causing the model to prefer lower weight values. The equation for L2 regularization is given by:
@@
where:
  • $\lambda$ is the regularization parameter (also called the penalty term or weight decay), which controls the strength of the regularization.
  • $w_{i}$ is the weight for the $i$-th feature in the model.
  • $n$ is the total number of weights in the model.

### v2

@@ Abstract
Open Radio Access Network (O-RAN) adopts a flexible, open, and virtualized architecture with standardized interfaces, reducing dependency on a single supplier. O-RAN hosts many intelligent applications known as xApps. xApps are applications deployed at the RAN Intelligent Controller (RIC) that leverage advanced AI/ML algorithms to make dynamic decisions for network optimization. Each application operates with distinct optimization objectives and is managed by independent operators while accessing shared network resources. Conflicts in this context occur when a deployed xApp’s objective interferes with another xApp, resulting in incompatible actions or decisions that may negatively impact network performance. The lack of a unified mechanism to coordinate and prioritize the actions of different applications can create three types of conflicts (direct, indirect, and implicit). Conflict management in O-RAN refers to the process of identifying and resolving conflicts between network applications. In our paper, we introduce a novel data-driven GCN-based method called GRAPH-based Intelligent xApp Conflict Prediction and Analysis (GRAPHICA) based on Graph Convolutional Network (GCN). It predicts three types of conflicts (direct, indirect, and implicit) and pinpoints the root causes (xApps). GRAPHICA captures the complex and hidden dependencies among the xApps, controlled parameters, and KPIs in O-RAN to predict possible conflicts. Then, it identifies the root causes (xApps) contributing to the predicted conflicts. The proposed method was tested on highly imbalanced synthesized datasets where conflict instances range from 40% to 10%. The model is tested in a setting that simulates real-world scenarios where conflicts are rare to assess its performance. Experimental results demonstrate a high F1-score over 98% for the synthesized datasets with different levels of class imbalance.
@@
5G Networks and Beyond, O-RAN, Quality of Service, Conflict Management, xApps, Graph Neural Network, Root Cause Analysis, Predictive Maintenance, Key Performance Indicators.
@@ I Introduction
Other studies seek to detect or predict the occurrences of conflicts in O-RAN. In [16], a team deep Q-learning algorithm for resource allocation is disclosed to mitigate xApps conflicts in O-RAN. In [17], a bargaining game-theoretic approach is introduced along with a standardized Conflict Mitigation System (CMS) and independent controllers. This work is further improved in [18]. The authors use a cooperative game theory to optimize the conflicting parameters while maximizing the number of xApps that meet QoS requirements. In [19], a Deep Q-Network and a multi-headed multilayer perceptron network model are presented. The method distills knowledge from multiple xApps to train a single model retaining the capabilities of previous xApps.
Other works address the xApps conflict management using graph-based approaches. The authors in [20] propose a method to detect, characterize, and mitigate conflicts among applications in O-RAN. Accordingly, they combine hierarchical graphs with statistical knowledge of applications to find dependencies between control parameters and Key Performance Measurements. In [21], the paper suggests using GraphSAGE for reconstructing and labeling conflict graphs in O-RAN. It captures the hidden dependencies between xApps, parameters, and KPIs. Then, graph labeling is performed to identify the different types of conflicts based on well-defined, graph-based definitions.
Understanding the complex and intricate dependencies between the xApps, the controlled parameters, and the KPIs is imperative to address the xApp conflict RCA. It allows us to predict the different types of conflicts in real time and derive the contributing root causes (xApps) for proactive maintenance. In the previously mentioned non-graph-based approaches, there is still a substantial limitation in their capacity to effectively capture and model dependencies within complex data structures. The proposed method in [21] leverages GraphSAGE to effectively capture the dependencies within the data. However, it still falls short of fully accounting for the imbalanced nature of the datasets in practical life scenarios, as these conflicts are rare. Addressing this limitation is crucial for enhancing the robustness and generalizability of the proposed approach in real-world applications. Accordingly, we propose GRAPHICA, a GCN-based method that is capable of handling highly imbalanced datasets with up to 10% conflicts. It predicts the xApps conflicts (direct, indirect, and implicit) and identifies their root causes (xApps) for future mitigation strategies.
@@
  • We propose a 5G O-RAN xApps conflict prediction with a root cause analysis method using Graph Convolutional Networks (GCNs). This allows us to capture the complex and hidden dependencies between the xApps, parameters, and KPIs in the network to predict xApps conflicts (direct, indirect, and implicit).
  • We leverage a binary state-based GCN structure to map and explain the complex relationships. This ensures the generalizability of the proposed solution, which can be extended to other networks. This structure along with the predictions are the key inputs for the root cause analysis.
  • We use the focal loss function in the model training. This allows the model to handle highly imbalanced datasets. The model maintains an F1 score greater than 98% for the synthesized datasets with different levels of class imbalance.
The remainder of this paper is structured as follows. Section II presents the problem formulation, clearly defining the research question and outlining the scope and objectives of the study. Section III describes the proposed system model. Section IV details the evaluation metrics, the experimental setup, and the interpretation of the results. Section  V discusses the results derived from the analysis. Section VI concludes this paper.
@@ II Preliminaries
Figure 1: Examples of xApps conflicts.
@@ III Proposed Method
A database is required to store the collected data from the SM in the Near-RT RIC using the messaging infrastructure. Then, the proposed system retrieves the collected data from the database for further processing using a shared data layer. Our proposed solution requires the implementation of a component that converts the collected data into a binary-state dataset, as depicted in Fig. 4. In particular, the state of an activated xApp is presented by “1”, and a deactivated xApp is represented by “0”. Similarly, if the value of a parameter or a KPI is changed from its previous timestamp value, the state is set to ”1”, otherwise, the state is set to ”0”.
The proposed solution executes three major tasks, namely Graph Structure Creator (GSC), Graph Anomaly Predictor (GAP), and Root Cause Analyst (RCA). Fig.  3 illustrates the chronological sequence of events, highlighting the progression from system monitoring to conflict occurrence. The predictive model utilizes real-time data and a GCN-based model to identify key triggers based on dynamic variables in the network and forecast the likelihood of a conflict occurrence. Specifically, at time t0, the proposed system initiates the monitoring process and undergoes a collection for the state dynamic behavior, including the xApps subscription, operational parameters, and KPIs values. At t1, an alert is raised based on the observed patterns if there is a likelihood of a deviation or anomaly in the system’s behavior and the possible root causes (xApps) are identified to trigger a need for immediate attention and intervention. If the alert is raised in time before the conflict occurs at t2, we can avoid a later KPI degradation at t3. In the remainder of this section, we will describe each module in the proposed method, outlining the corresponding steps and functionality.
@@
Figure 3: The chronological sequence of events for conflicts prediction.
III-A Data Curation and Conflicts Modeling
@@ III-B Graph Structure Creator (GSC)
The GSC module is responsible for creating graph-structured data from the binary-state dataset comprising the collected data from the Near-RT RIC (see Fig. 5). The GCS ingests the binary-state dataset and the number of xApps, parameters, and KPIs as input and convert them into graph-structured data for further processes using three subgraphs: $G^{\mathrm{PA}}$, $G^{\mathrm{KP}}$, and $G^{\mathrm{P^{\prime}P}}$.
Figure 5: The graph structure creator (GSC) module.
@@
The subgraph $G^{\mathrm{PA}}=\left(V^{\mathrm{PA}},E^{\mathrm{PA}}\right)$ represents the relationships between the applications and the control parameters. Nodes of $G^{\mathrm{PA}}$ are both the applications and the control parameters, i.e., $V^{\mathrm{PA}}=\mathcal{P}\cup\mathcal{A}$, and edges $E^{\mathrm{PA}}$ represent whether or not an application is controlling a parameter $p\in\mathcal{P}$. Any 2-tuple $(p,a)\in\mathcal{P}\times\mathcal{A}$ is an edge of $G^{\mathrm{PA}}$, i.e., $(p,a)\in E^{\mathrm{PA}}$ if and only if parameter $p$ is directly controlled by an application $a$. This will cause a direct conflict.
@@
The subgraph $G^{\mathrm{KP}}=\left(V^{\mathrm{KP}},E^{\mathrm{KP}}\right)$ represents the relationships between the control parameters and the KPIs. Nodes of $G^{\mathrm{KP}}$ are both the parameters and the KPIs, i.e., $V^{\mathrm{KP}}=\mathcal{K}\cup\mathcal{P}$, and edges $E^{\mathrm{KP}}$ represent whether or not a parameter is modifying the value of a KPI $p\in\mathcal{P}$. Any 2-tuple $(k,p)\in\mathcal{K}\times\mathcal{P}$ is an edge of $G^{\mathrm{KP}}$, i.e., $(k,p)\in E^{\mathrm{KP}}$ if and only if KPI $k$ is directly affected by a parameter $p$. This will cause an indirect conflict.
@@
The subgraph $G^{\mathrm{P^{\prime}P}}=\left(V^{\mathrm{P^{\prime}P}},E^{\mathrm{P^{\prime}P}}\right)$ represents the relationships between the control parameters themselves. Nodes of $G^{\mathrm{P^{\prime}P}}$ are both the parameters in the sets $\mathcal{P^{\prime}}$ and $\mathcal{P}$, i.e., $V^{\mathrm{P^{\prime}P}}=\mathcal{P^{\prime}}\cup\mathcal{P}$, and edges $E^{\mathrm{P^{\prime}P}}$ represent whether or not a parameter $p\in\mathcal{P}$ is modifying the value of another parameter $p^{\prime}\in\mathcal{P^{\prime}}$. Any 2-tuple $(p^{\prime},p)\in\mathcal{P^{\prime}}\times\mathcal{P}$ is an edge of $G^{\mathrm{P^{\prime}P}}$, i.e., $(p^{\prime},p)\in E^{\mathrm{P^{\prime}P}}$ if and only if a parameter $p^{\prime}$ is directly affected by a parameter $p$. This will cause an implicit conflict.
@@
To construct the graph $G^{\mathrm{PA}}$, the GSC iterates over the applications and the parameters nodes and stores their corresponding state values. Subsequently, the resultant graph nodes are xApps and the parameters. The edges represent the nodes having a state “1” concurrently. The edges reflect the nodes that are affecting each other at the same time. This leads to direct conflicts.
@@
Similarly, for $G^{\mathrm{KP}}$, the GSC iterates over the parameters and the KPIs nodes and stores their corresponding state values. The output graph nodes are the parameters and the KPIs. The edges indicate the parameters and KPIs whose values change simultaneously by with a state “1”, resulting in implicit conflicts.
@@
Lastly, for $G^{\mathrm{P^{\prime}P}}$, the GSC iterates over the parameters and checks their state values. The nodes of the graphs are the parameters, and the edges refer to the links among the parameters that have a state of “1”. This means the parameters are impacting each other’s behavior. This produces indirect conflicts.
@@
During the construction of the three subgraphs, unique node features and edge attributes are assigned to the nodes and their edges, contributing to the same type of conflict. GCN uses node features as the starting point to update the node embeddings (representations) through multiple layers of convolutions. These embeddings are refined iteratively as the model propagates information across neighboring nodes. Including edge attributes in the graph convolution process allows the model to learn a more nuanced understanding of node interactions. When performing convolution, edge attributes help the GCN understand the nature of the relationship between nodes.
III-C Graph Anomaly Predictor (GAP)
@@
The GAP module is responsible for predicting whether the instance is normal or whether we have a conflict (see Fig. 6). The output of GAP will be one of four labels (0: normal, 1: direct conflict, 2: implicit conflict, and 3: indirect conflict). The fully constructed graph structure from the GSC module is fed to the GAP module. It consists of a Graph Convolutional Network (GCN). GCN learns node-level representations from the graph-structured data, which are then aggregated via a global mean pooling into a graph-level representation for final classification. Then, a fully connected neural network (NN) layer is used to learn the final transformation from the graph’s representation to the output classes.
Figure 6: The graph anomaly predictor module.
@@
To address the imbalances in the datasets, the classifier is enhanced by a focal loss function [24]. It improves the standard cross-entropy loss by adding two key parameters, alpha and gamma. This function adjusts the loss contribution of easy and hard examples. It focuses more on hard-to-classify examples that are misclassified by the model. In parallel, it down-weights the loss associated with easy-to-classify examples. $\alpha_{c}$ adjusts for class imbalance by balancing the weight between different classes. $\gamma$ focuses the model on harder examples by reducing the loss for easy, well-classified examples. This helps to prevent the model from being overwhelmed by the majority class in imbalanced datasets and gives more attention to challenging instances. We compute $\alpha_{c}$ using inverse class weights by calculating the inverse frequency of each class in the dataset. On the other hand, $\gamma$ is manually tuned for model optimization. The focal loss function for multi-class classification is given by:
@@
Instead of processing the entire graph at once, mini-batch processing is executed. This involves processing smaller subgraphs simultaneously, which helps reduce memory usage and computational time. Furthermore, a stratified K-Fold is implemented to ensure that each fold in the cross-validation process has a similar distribution of classes as the entire dataset during training and validation. This is particularly useful since real-world traffic is highly imbalanced. Early stopping is implemented in our model which is a method used to prevent overfitting by monitoring the validation loss during training, and halting the training process if the model’s performance on the validation set does not improve for a specified number of epochs. The delta ensures that only considerable improvements in validation loss are evaluated as indicators of progress, preventing early stopping from being triggered by negligible changes.
The L2 regularization is implemented to mitigate overfitting. The L2 regularization adds a penalty to the loss function that prevents the model from giving too much weight to any single feature. It achieves this by squaring the model’s weights and adding the result to the loss, causing the model to prefer lower weight values. The equation for L2 regularization is given by:
@@
where $\lambda$ is the regularization parameter controlling the strength of the regularization, $w_{i}$ is the weight for the $i$-th feature in the model, and $n$ is the total number of weights in the model.